tan of phata if the sin of phata is the square root of 7 divided by 6
Answer by lwsshak3(6505)

tan of phata if the sin of phata is the square root of 7 divided by 6
**
use x for phata
given: sinx=√7/6=opposite side/hypotenuse
adjacent side=√[6^2-(√7)^2]=√(36-7)=√29
tanx=opp side/adj side=√7/√26